OpenStudy (anonymous):

k=1 sigma k tends to infinity (-3/2)^(k+1) determine whether series converges or diverges if series is converges find the sum

OpenStudy (anonymous):

hey did you get my answer to the last one? the site froze and i could not complete it

OpenStudy (turingtest):

I gave an answer to this already

OpenStudy (turingtest):

\[\sum_{k=1}^\infty(-\frac32)^2(-\frac32)^{k-1}\]

OpenStudy (experimentx):

did you mean?? $$ \sum_{k = 1}^{\infty} \left ( \frac{-3}{2} \right ) ^{k+1}$$ also don't close question and ask new one. redundancy is no better than spam

OpenStudy (anonymous):

not much to this one as the terms grow infinitely large for k odd (k+1 even) and go to minus infinity if k is even (so k + 1 is odd) cannot converge

OpenStudy (anonymous):

turing test u not clearly ans this question half ans u give me not a final answer

OpenStudy (experimentx):

The converging alternate seris should be like 1) and the diverging alternate series whousl be like 2)